Question 2, chap 123, sect 5. part 2 of 2 10 points What is the angle 0 between the electro- static force on the charge at the origin and the positive x-axis? Answer as an angle between 180° and 180° measured from the positive r-axis, with counterclockwise positive Answer in units of Question 3, chap 123, sect 5. part 1 of1 10 points Consider the three charges below The value of Coulombs constant is 8.9875 x 10 N.m2 /c2 3 HC 90° 9 HiC What is the magnitude of the electrostatic force |F|| on the top charge? Answer in units ofN Show step by step solution for both questions please
Answer #1

question 3

==========

force due to 9uC

F1 = k q1 q / r^2

F1 = 9*10^9* 9* 3*10^-12/ 0.49^2 = 1.01 N

force due to - 9 uC

F2 = 9*10^9* 9*3*10^-12 / 0.49^2 = 1.01 N

since both the force acts at right angle, so net force is given by

F^2 = F1^2 + F2^2

F = 1.431 N

=========

Problem 2 : part (1) also needed.

=========

do comment in case any doubt, will reply for sure.. Goodluck